# REGULAR ARTICLE

### v1

@@ 5 The way from here
— Jonathan P. Dowling [78]
Throughout this article, we have outlined a trajectory from 19th-century thermodynamics to modern formalisms, passing through the emergence of stochastic thermodynamics and culminating in the rise of quantum thermodynamics [79], formulated here within the language of LRT near equilibrium. In this framework, the QFT for TPM-based dissipated work reaches a height without relying on slow-driving or weak coupling aproximations [57]. This perspective may shed light on non-Markovian effects and work fluctuations across phase transitions [80], while remaining experimentally relevant since the Fano factor is directly measurable [57].
While the TPM definition is useful, quantum work is not an observable [72], and the approach suffers from serious drawbacks [21]. In particular, it neglects measurement back-action, and projective energy measurements are neither feasible nor practical [21]. These limitations motivate describing work fluctuations in terms of variations of physical quantities (VPQs) [81]. A promising alternative is the two-time observable protocol, which is experimentally accessible in platforms such as trapped ions, plasma drops, superconducting qubits, and NMR setups [81]. In this context, linear response theory may offer a natural framework for modeling these systems.
As our knowledge advances, we are steadily moving toward the ability to design highly efficient, optimised devices—such as quantum computers, sensors, and metrological tools — that harness quantum phenomena as fundamental resources. This heralds the onset of the second quantum revolution [78], in which quantum effects extend beyond atoms and molecules to profoundly shape technology, society, and civilisation.
6 Acknowledgements

### v2

@@ V The way from here
— Jonathan P. Dowling dowling2003quantum
Throughout this article, we have outlined a trajectory from 19th-century thermodynamics to modern formalisms, passing through the emergence of stochastic thermodynamics and culminating in the rise of quantum thermodynamics halpern2022quantum , formulated here within the language of LRT near equilibrium. In this framework, the QFT for TPM-based dissipated work reaches a height without relying on slow-driving or weak coupling aproximations guarnieri2024generalized . This perspective may shed light on non-Markovian effects and work fluctuations across phase transitions campbell2025roadmap , while remaining experimentally relevant since the Fano factor is directly measurable guarnieri2024generalized .
While the TPM definition is useful, quantum work is not an observable talkner2007fluctuation , and the approach suffers from serious drawbacks deffner2019quantumthermo . In particular, it neglects measurement back-action, and projective energy measurements are neither feasible nor practical deffner2019quantumthermo . These limitations motivate describing work fluctuations in terms of variations of physical quantities (VPQs) silva2025fundamental . A promising alternative is the two-time observable protocol, which is experimentally accessible in platforms such as trapped ions, plasma drops, superconducting qubits, and NMR setups silva2025fundamental . In this context, linear response theory may offer a natural framework for modeling these systems.
As our knowledge advances, we are steadily moving toward the ability to design highly efficient, optimised devices—such as quantum computers, sensors, and metrological tools — that harness quantum phenomena as fundamental resources. This heralds the onset of the second quantum revolution dowling2003quantum , in which quantum effects extend beyond atoms and molecules to profoundly shape technology, society, and civilisation.
VI Acknowledgements